# Score-based diffusion nowcasting of GOES imagery

### v1

@@ 3.2.1 U-Net architecture
The U-Net choice follows an online tutorial available on HuggingFace (archived into the github repository for this paper, see Data Availability). It is a fairly large U-Net (Ronneberger et al., 2015) with six up- and down- sampling blocks, which contain two res-net blocks (He et al., 2015) each. There is also attention layers added at the fifth of the six total blocks and inside the bottleneck. This model has 113 million learnable parameters. We acknowledge that only using one core architecture limits the generalization of this paper, but due to long training times (order of days) on our GH-200, we are limited in our ability to perform hyperparameter tuning of the architecture. Using smaller core models (i.e., with fewer parameters) and core models without attention will likely speed up training and inference times for satellite nowcasting and is a topic of future research.
@@ 3.2.3 Diffusion models
We train three different diffusion variants that all use the U-Net as denoiser. The first diffusion model, named Diff follows the Karras et al. (2022) framework, where the denoiser is denoising the 10 min forecast. The second diffusion model, named CorrDiff, follows the approach from Mardani et al. (2025) and Pathak et al. (2024). It uses the pre-trained U-Net of the baseline model to get an initial, potentially blurry, prediction of the 10 min forecast. Then it trains a denoiser to improve the forecast of the plain U-Net. The denoiser uses the same condition as Diff, along with the estimate obtained from the pre-trained U-Net output (i.e., the input is now 4 channels). The other change compared to Diff is that the CorrDiff model predicts the difference between the plain U-Net estimate and the truth.
@@
Given the relatively long training (evaluation) time for the Diff and CorrDiff models, we also explore a latent diffusion model (e.g., Leinonen et al., 2023), named LDM. That is, we use a pre-trained variational autoencoder (VAE) from HuggingFace to compress our original data from 256 by 256 (with one channel) down to 64 by 64 pixels (with four channels). Due to the channel dimension going up, the compression factor is only four. The choice of VAE is discussed in Section 33.3, but know that we choose the autoencoder that has the smallest reconstruction loss on the GOES imagery (Table 1). We then use the compressed data to train the denoiser for diffusion in this latent space. This results in a five-times speed-up in training, where the latent model completed training for 1,000 epochs in about 1 day. We did attempt building a latent version of the CorrDiff model, but the U-Net predictions in the latent space were poor on their own and we decided to not pursue it further for this paper.
@@ 3.3 Off-the-shelf Variational Auto Encoders
The latent diffusion model requires the use of a variational autoencoder. Here we discuss the use of pre-trained variational autoencoders (VAE) for this purpose, specifically for compressing meteorological imagery. The availability of open-source pre-trained models have accelerated many efforts in the computer science field (e.g., Stable Diffusion; Rombach et al., 2022). Given these pre-trained autoencoders are trained on pictures of cats, dogs, people, places and things, we were curious if these pre-trained models could be used accurately on meteorological satellite data. We tested eight total VAEs and the reconstruction losses for each one are found in Table 1. In general, most of them have relatively poor reconstructions, showing RMSEs of two and three kelvin where any individual pixel could be upwards of 20-30 K different than the original image. There is one that we tested that did have reconstruction losses less than 1 K. This VAE (example shown in Fig. 7) was trained to perform super-resolution, and encodes the data spatially. Given the low reconstruction loss, we use this VAE for compressing our original satellite data from 256 by 256 pixels to 64 by 64 and training the diffusion in the 64 by 64 space (latent diffusion model). As will be seen later, while the latent diffusion can be skillful, it has some undesired qualities that might be a result of the VAE itself. Tuning the pre-trained VAE or training a VAE from scratch on satellite data specifically are both topics of future research.
@@ 3.4 Generation details and ensembles
As described in Section 2, diffusion machine learning tasks are split into training the denoiser and generation. With this split come a few additional hyperparameters solely linked to the generation procedure. These generation hyperparameters are centered on: how many denoising steps to take (num_steps); the maximum noise level to start from (sigma_max); the spacing of the steps (rho); the amount of stochastic noise injected into the denoising process (S_churn); and the scale of noise added back in (S_noise). Just like training hyperparameters, a hyperparameter search would ideally be done for all diffusion models trained. Unfortunately, given the computational demand to run Diff and the CorrDiff model, we were only able to do a sparse hyperparameter search on the latent diffusion model and assume that these parameters work well for the others. This assumption may be false, leading to suboptimal parameter sets, but this is an area of future research. More specifically, we perform a gridsearch of four of the above hyperparameters for the latent model and then choose the best joint performance of root mean squared error and the spread–skill ratio on the 1,000 example validation (i.e., use 10 different noisy images to start with the same two condition images. The specific values in the gridsearch are in Table 2.
@@
Notice the mention of the spread-skill ratio, which is a metric for ensemble skill (Haynes et al., 2023). A benefit of the diffusion based models is the ability to generate as many samples as desired. This is partly controlled by the hyperparamters discussed above but also by varying the choice noisy image along the zeroth dimension of the input data to the neural network (see Fig. 5 and Fig. 6b). For the ensembles analyzed here, we vary this noisy dimension with 10 different noise seeds to then get 10 different realizations of the same forecast. This is done for all three diffusion models. We choose 10 here as a trade off between the memory (time) required for running the forecast and having a decent ensemble size. For the Diff and the CorrDiff models it takes approximately 10 minutes to generate 10 members of a three hour forecast on a single NVIDIA GH-200. LDM takes about one minute to perform the same task.
@@ 3.5 Autoregression
All machine learning forecast methods trained here follow the structure of many of the data-driven weather prediction models and optical flow techniques where two images ($\tau_{0}$, $\tau_{-10}$) are used to forecast a single time step ($\tau_{+10})$ (Fig. 6b and Fig. 8). Then, the output from that single forecast can be fed back into the forecast models recursively. Forecasts of arbitrary length can then be rolled out. We choose to forecast out three hours (18 autoregressive steps). Some of the data driven models initially train on the single step prediction paradigm but then tune with a low learning rate on multiple steps. It is not clear how that could be done with score-based diffusion since the training of the neural network is separated from the image generation but this is an area of future research. Thus, no multi-step fine tuning is done.
@@ 3.6 Metrics
4 Results and Discussion
We first present results for a case study to provide visual intuition for each model’s characteristics, followed by performance metrics for the entire test set.
@@ 4.1 Case Study: The coast of South America
We provide an in depth case study that has a variety of cloud types to evaluate how well the various forecasts work with the gambit of clouds observed in the atmosphere. The case we show was chosen out of the validation set such that it contained low level and high level clouds, as well as clouds that form and decay over the three hour forecast. An overview of the case is shown in Fig. 9 showing the visible reflectance, the cloud top height retrieved from NOAA’s Enterprise AWG Cloud Height Algorithm (ACHA; Heidinger et al. (2020)), and the observed infrared brightness temperatures.
@@
For the Diff approach, which uses the plain diffusion model in Karras et al. (2022)), the forecast loosely evolves like the observations. The stratocumulus deck recedes to the west, and has pieces of it break up. A surprising result was the ability of the diffusion model to initiate convection close to where convection was observed despite only knowing about 20 mins of observed satellite data. As a reminder, all models here have no information on the surface type, where the sun is, what time of day it is, etc. It is remarkable that the diffusion model can evolve a realistic looking scenario off of such limited information. Not all members of the diffusion model have convection in the same spot, evidenced by the ensemble mean being a bit smoother and no clear cold clouds like in the single member scenario. Clearly, the diffusion model is not perfect. The overall scene seems a bit noisy, suggesting the hyperparameters of the generation of forecasts are suboptimal. Furthermore, there is an overal warm bias of the scene, where the brightness temperature in the forecast are warmer than truth (blue hues in the third row of Fig. 11). Overall we are encouraged by the ability of the diffusion model to evolve the scene and looking realistic (i.e., not blurry) in doing so out to three hours.
@@
For the CorrDiff approach, which uses the plain U-Net forecast as additional input, we can see that the scenes are considerably less noisy than for the Diff model (i.e., looks clearer). It could be that the hyperparameter set of the generation procedure is more optimal, or that the U-Net provides a good initial guess of the state and the diffusion can then just focus on small scale features. Mardani et al. (2025) suggest that the use of the U-Net helps reduce the variance of the task and makes it easier for the machine learning to learn. The CorrDiff model also convects the atmosphere like Diff model, but this time the convection shows up further to the NE than the observed convection. The stratocumulus retreat to the west is much better resolved by the CorrDiff model and does have a few breaks in the cloud deck. Also, as in the other forecast models, the cirrus streamer is not well captured beyond an hour, likely because of its tenuous nature. Similarly to the Diff model, the ensemble mean does not have a single cold cloud feature like the single member, but does have an area of colder clouds to the NE where cumulus and convection were observed. The CorrDiff model is not immune to the drift of the other forecast and infact overcorrects the U-net input to now have a cold bias (red hues in the fourth row of Fig. 11).
@@
Considering the fastest of the diffusion models, the latent diffusion model (LDM), we can see for the first hour the evolution looks similar to the other diffusion models. After an hour, the edges of the image have a strogn cold bias (red hues in the fifth row of Fig. 11). We suspect this is because of the non-optimal encoding of the satellite images with the pre-trained autoencoder. If you consider Fig. 7b, the edge of the image has some odd cold/hot pixels. This was common when inspecting the latent space (not shown) and leads to large errors in the encoding-decoding process (Fig. 7d). The evolution of the surface warming also looks off compared to the other diffusion models and the observed data. Thus, while the LDM model can work, showing an evolution of clouds and even initiating convection, there are clear tradeoffs for its ability to run five to ten times faster. Furthermore, as mentioned before, the VAE used here is likely not yet optimal.
@@
Figure 12: A measure of image sharpness, fractional powerspectra change for the images in (Fig. 10). The light blue line is the Diff forecast, purple line is CorrDiff, dark blue is the LDM forecast and the plain U-net is in orange. The dashed lines are the ensemble means for each of the corresponding forecasts. (a) the 10 min forecase (b) the 3 hour forecast.
@@
A motivation of this paper was to enhance the often blurry output of neural networks that optimize some variant of the MSE loss. To quantify this a bit further, we consider the power spectra of each nowcast for the same case and compare it to the original image. Fig. 12 shows the fractional change for the 10 min forecast and the three hour forecast (Fig. 10). For the 10 min forecast, for wavelengths greater than 10 km all match well. For smaller features, the single member diffusion (solid light blue) and single member CorrDiff (solid purple) have similar or more of the original features than the U-Net and the ensemble means (dashed lines). For waves smaller than 4 km (two times the native resolution) the U-Net effectively removes these features despite the image in Fig. 10 looking fairly sharp. Similarly the ensemble means reduce the number of small waves (less than five kilometers), but not to the extent of the U-Net and this smoothing is expected given there could be divergence among the ensemble members. Surprisingly, the LDM (dark blue line) and to a less extent the Diff model (light blue line) add small scale features. This is likely a result of a leftover noise from the denoising process of diffusion.
@@
At the three hour forecast, the spectra are a bit harder to interpret because of how different the forecasts are from one another. Clearly, the Diff model has more waves at all wavelengths, which is the noisy feature we described in Fig. 10. The ensemble means are all smoother than the U-Net, expected from the spread of the ensemble members. Even the CorrDiff model (solid purple line) shows some blurriness overall compared to the observed image, but does have larger spectra for waves smaller than four km, compared to the U-Net. For future research we plan to use additional sharpness metrics to better quantify the impact of diffusion on overall sharpness of the machine learning output (Ebert-Uphoff et al., 2025).
@@ 4.2 Bulk results
We compare here mainly to persistence and the U-Net forecast. Fig. 13 confirms what was shown qualitatively in Fig. 10 where most of the diffusion models have added value over the more traditional U-Net approach and a simple persistence forecast. Furthermore, all methods not only advect clouds (e.g., an optical flow solution; Shakya and Kumar, 2019), they can also create and decay clouds. We are encouraged that these models work given the limited information content included as inputs.
@@
A key perk that we have seen already in the diffusion framework is the ability to generate n-many samples to use ensemble based forecasts. This already results in improved pixel-based metrics in Fig 13. We can also look to quantify how well the uncertainty of that ensemble corresponds to the true error of the model. We leverage the skill-spread diagram (e.g., Haynes et al., 2023; Delle Monache et al., 2013) where the standard deviation of the 10 member diffusion ensembles is compared to the RMSE of the ensemble (Fig. 14).
@@
5 Summary and Conclusions
A challenge with many machine learning methods is the prevalence of blurry solutions that then have limited usefulness in human-in-the-loop applications (i.e., weather forecasters). Diffusion models in the literature suggest to be a solution to these blurry outputs while also providing useful uncertainty quantification through ensembles. Thus, for this paper we explored the use of score-based diffusion models to nowcast geostationary satellite imagery (i.e., GOES infrared brightness temperature). The nowcast was built with the eventual use-case of forecasting cloud and precipitation location so evaluations were centered on the evolution of cold brightness temperature regions in the imagery.
The first contribution of this work was a deep dive into the methodology of score-based diffusion models (Section 2). In the deep dive, we discussed some of the math as well as some base intuition behind score-based diffusion models. Furthermore, we discussed how to train and use score-based diffusion models in an attempt to accelerate their use in the Earth and Environmental Sciences.
The second contribution of this work was the exploration of three different types of score-based diffusion models and how well they work for a simplified nowcasting task of satellite imagery. More specifically, we explore the use of three score-based diffusion models, a diffusion model (Diff) that follows Karras et al. (2022), a residual corrective diffusion model (CorrDiff) that follows Mardani et al. (2025) and a latent diffusion model (LDM) inspired by Leinonen et al. (2023) to forecast GOES channel 13 infrared brightness temperatures out to three hours. We showed that the diffusion models were able to nowcast satellite imagery qualitatively and quantitatively better than both a persistence forecast and a mean squared error trained U-Net (Figures 10-14). The best performing diffusion model was the CorrDiff approach followed by the Diff model and then the LDM. All diffusion models were able to not only advect the current clouds in the scene but also generate and decay clouds. We also showed that all methods were able to generate convection (i.e., localized regions of very cold brightness temperatures) despite only having two images (20 mins) of 10 micron infrared imagery and no information on the thermodynamic state of the atmosphere, the solar input, or the land surface type.

### v2

@@ 3.2.1 U-Net architecture
The U-Net choice follows an online tutorial available on HuggingFace (archived into the github repository for this paper, see Data Availability). It is a fairly large U-Net (Ronneberger2015) with six up- and down- sampling blocks, which contain two res-net blocks (He2015) each. There is also attention layers added at the fifth of the six total blocks and inside the bottleneck. This model has 113 million learnable parameters. We acknowledge that only using one core architecture limits the generalization of this paper, but due to long training times (order of days) on our GH-200, we are limited in our ability to perform hyperparameter tuning of the architecture. Using smaller core models (i.e., with fewer parameters) and core models without attention will likely speed up training and inference times for satellite nowcasting and is a topic of future research.
@@ 3.2.3 Diffusion models
We train three different diffusion variants that all use the U-Net as denoiser. The first diffusion model, named Diff follows the Karras2022 framework, where the denoiser is denoising the 10 min forecast. The second diffusion model, named CorrDiff, follows the approach from Mardani2025 and Pathak2024. It uses the pre-trained U-Net of the baseline model to get an initial, potentially blurry, prediction of the 10 min forecast. Then it trains a denoiser to improve the forecast of the plain U-Net. The denoiser uses the same condition as Diff, along with the estimate obtained from the pre-trained U-Net output (i.e., the input is now 4 channels). The other change compared to Diff is that the CorrDiff model predicts the difference between the plain U-Net estimate and the truth.
@@
Given the relatively long training (evaluation) time for the Diff and CorrDiff models, we also explore a latent diffusion model (e.g., Leinonen2023), named LDM. That is, we use a pre-trained variational autoencoder (VAE) from HuggingFace to compress our original data from 256 by 256 (with one channel) down to 64 by 64 pixels (with four channels). Due to the channel dimension going up, the compression factor is only four. The choice of VAE is discussed in Section 33.3, but know that we choose the autoencoder that has the smallest reconstruction loss on the GOES imagery (Table 1). We then use the compressed data to train the denoiser for diffusion in this latent space. This results in a five-times speed-up in training, where the latent model completed training for 1,000 epochs in about 1 day. We attempted a latent CorrDiff model by training the corresponding U-Net model entirely in the latent space, but the U-Net predictions in the latent space were poor on their own and we decided to not pursue it further for this paper.
@@ 3.3 Off-the-Shelf Variational Auto Encoders
The latent diffusion model requires the use of a variational autoencoder. Here we discuss the use of pre-trained variational autoencoders (VAE) for this purpose, specifically for compressing meteorological imagery. The availability of open-source pre-trained models have accelerated many efforts in the computer science field (e.g., Stable Diffusion; Rombach2022). Given these pre-trained autoencoders are trained on pictures of cats, dogs, people, places and things, we were curious if these pre-trained models could be used accurately on meteorological satellite data. We tested eight total VAEs and the reconstruction losses for each one are found in Table 1. In general, most of them have relatively poor reconstructions, showing RMSEs of two and three kelvin where any individual pixel could be upwards of 20-30 K different than the original image. There is one that we tested that did have reconstruction losses less than 1 K. This VAE (example shown in Fig. 9) was trained to perform super-resolution, and encodes the data spatially. Given the low reconstruction loss, we use this VAE for compressing our original satellite data from 256 by 256 pixels to 64 by 64 and training the diffusion in the 64 by 64 space (latent diffusion model). As will be seen later, while the latent diffusion can be skillful, it has some undesired qualities that might be a result of the VAE itself. Tuning the pre-trained VAE or training a VAE from scratch on satellite data specifically are both topics of future research.
@@ 3.4 Generation Details and Ensembles
As described in Section 2, diffusion machine learning tasks are split into training the denoiser and generation. With this split come a few additional hyperparameters solely linked to the generation procedure. These generation hyperparameters are centered on: how many denoising steps to take (num_steps); the maximum noise level to start from (sigma_max); the spacing of the steps (rho); the amount of stochastic noise injected into the denoising process (S_churn); and the scale of noise added back in (S_noise). Just like training hyperparameters, a hyperparameter search would ideally be done for all diffusion models trained. Unfortunately, given the computational demand to run Diff and the CorrDiff model, we were only able to do a sparse hyperparameter search on the latent diffusion model and assume that these parameters work well for the others. This assumption may be false, leading to suboptimal parameter sets, but this is an area of future research. More specifically, we perform a gridsearch of four of the above hyperparameters for the latent model and then choose the best joint performance of root mean squared error and the spread–skill ratio on the 1,000 example validation (i.e., use 10 different noisy images to start with the same two condition images). The specific values in the gridsearch are in Table 2.
@@
Notice the mention of the spread-skill ratio, which is a metric for ensemble skill (Haynes2023). A benefit of the diffusion based models is the ability to generate as many samples as desired. This is partly controlled by the hyperparamters discussed above but also by varying the choice noisy image along the zeroth dimension of the input data to the neural network (see Fig. 7 and Fig. 8b). For the ensembles analyzed here, we vary this noisy dimension with 10 different noise seeds to then get 10 different realizations of the same forecast. This is done for all three diffusion models. We choose 10 here as a trade off between the memory (time) required for running the forecast and having a decent ensemble size. For the Diff and the CorrDiff models it takes approximately 10 minutes to generate 10 members of a three hour forecast on a single NVIDIA GH-200. LDM takes about one minute to perform the same task.
@@ 3.5 Autoregression
All machine learning forecast methods trained here follow the structure of many of the data-driven weather prediction models and optical flow techniques where two images ($\tau_{0}$, $\tau_{-10}$) are used to forecast a single time step ($\tau_{+10})$ (Fig. 8b and Fig. 10). Then, the output from that single forecast can be fed back into the forecast models recursively. Forecasts of arbitrary length can then be rolled out. We choose to forecast out three hours (18 autoregressive steps). Some of the data driven models initially train on the single step prediction paradigm but then tune with a low learning rate on multiple steps. While technically possible to train diffusion through multiple inference steps, it would require very large amounts of VRAM not capable with a single GH-200.
@@ 3.6 Metrics
4 Results
We first present results for a case study to provide visual intuition for each model’s characteristics, followed by performance metrics for the entire test set. The nowcast was built with the eventual use-case of forecasting cloud and precipitation location, so our evaluations focus on the evolution of cold brightness temperature regions in the imagery.
@@ 4.1 Case Study: The Coast of South America
We provide an in depth case study that has a variety of cloud types to evaluate how well the various forecasts work with the gambit of clouds observed in the atmosphere. The case we show was chosen out of the validation set such that it contained low level and high level clouds, as well as clouds that form and decay over the three hour forecast. An overview of the case is shown in Fig. 11 showing the visible reflectance, the cloud top height retrieved from NOAA’s Enterprise AWG Cloud Height Algorithm (ACHA; heidinger2018enterprise), and the observed infrared brightness temperatures.
@@
For the Diff approach, which uses the plain diffusion model in Karras2022), the forecast loosely evolves like the observations. The stratocumulus deck recedes to the west, and has pieces of it break up. A surprising result was the ability of the diffusion model to initiate convection close to where convection was observed despite only knowing about 20 mins of observed satellite data. As a reminder, all models here have no information on the surface type, where the sun is, what time of day it is, etc. It is remarkable that the diffusion model can evolve a realistic looking scenario off of such limited information. Not all members of the diffusion model have convection in the same spot, evidenced by the ensemble mean being a bit smoother and no clear cold clouds like in the single member scenario. Clearly, the diffusion model is not perfect. The overall scene seems a bit noisy, suggesting the hyperparameters of the generation of forecasts are suboptimal. Furthermore, there is an overal warm bias of the scene, where the brightness temperature in the forecast are warmer than truth (blue hues in the third row of Fig. 13). Overall we are encouraged by the ability of the diffusion model to evolve the scene and looking realistic (i.e., not blurry) in doing so out to three hours.
@@
For the CorrDiff approach, which uses the plain U-Net forecast as additional input, we can see that the scenes are considerably less noisy than for the Diff model (i.e., looks clearer). It could be that the hyperparameter set of the generation procedure is more optimal, or that the U-Net provides a good initial guess of the state and the diffusion can then just focus on small scale features. Mardani2025 suggest that the use of the U-Net helps reduce the variance of the task and makes it easier for the machine learning to learn. The CorrDiff model also convects the atmosphere like Diff model, but this time the convection shows up further to the NE than the observed convection. The stratocumulus retreat to the west is much better resolved by the CorrDiff model and does have a few breaks in the cloud deck. Also, as in the other forecast models, the cirrus streamer is not well captured beyond an hour, likely because of its tenuous nature. Similarly to the Diff model, the ensemble mean does not have a single cold cloud feature like the single member, but does have an area of colder clouds to the NE where cumulus and convection were observed. The CorrDiff model is not immune to the drift of the other forecast and infact overcorrects the U-Net input to now have a cold bias (red hues in the fourth row of Fig. 13).
@@
Considering the fastest of the diffusion models, the latent diffusion model (LDM), we can see for the first hour the evolution looks similar to the other diffusion models. After an hour, the edges of the image have a strong cold bias (red hues in the fifth row of Fig. 13). We suspect this is because of the non-optimal encoding of the satellite images with the pre-trained autoencoder. If you consider Fig. 9b, the edge of the image has some odd cold/hot pixels. This was common when inspecting the latent space (not shown) and leads to large errors in the encoding-decoding process (Fig. 9d). The evolution of the surface warming also looks off compared to the other diffusion models and the observed data. Thus, while the LDM model can work, showing an evolution of clouds and even initiating convection, there are clear tradeoffs for its ability to run five to ten times faster. Furthermore, as mentioned before, the VAE used here is likely not yet optimal.
@@
Figure 14: A measure of image sharpness, fractional powerspectra change for the images in (Fig. 12). The light blue line is the Diff forecast, purple line is CorrDiff, dark blue is the LDM forecast and the plain U-Net is in orange. The dashed lines are the ensemble means for each of the corresponding forecasts. (a) the 10 min forecast (b) the 3 hour forecast.
@@
A motivation of this paper was to enhance the often blurry output of neural networks that optimize some variant of the MSE loss. To quantify this a bit further, we consider the power spectra of each nowcast for the same case and compare it to the original image. Fig. 14 shows the fractional change for the 10 min forecast and the three hour forecast (Fig. 12). For the 10 min forecast, for wavelengths greater than 10 km all match well. For smaller features, the single member diffusion (solid light blue) and single member CorrDiff (solid purple) have similar or more of the original features than the U-Net and the ensemble means (dashed lines). While the usefulness of evaluating the sharpness of the ensemble mean is questionable, we include it here to provide a comparison point for how blurry the individual members could get. For waves smaller than 4 km (two times the native resolution) the U-Net effectively removes these features despite the image in Fig. 12 looking fairly sharp. Similarly the ensemble means reduce the number of small waves (less than five kilometers), but not to the extent of the U-Net and this smoothing is expected given there could be divergence among the ensemble members. Surprisingly, the LDM (dark blue line) and to a less extent the Diff model (light blue line) add small scale features undetectable to the human eye in Fig. 12 but more noticeable in the the difference plot with red and blue pixels being located next to each other across the image (13). These small wave artifacts in the LDM model are from the VAE itself, which was found to add noise in the decoding process (not shown). Meanwhile, we hypothesize that the small wave artifacts in the Diff model are left over from the denoising process of the diffusion.
@@
At the three hour forecast, the spectra are a bit harder to interpret because of how different the forecasts are from one another but the changes to the smoothness in the imagery is clear. The Diff model has more waves at all wavelengths, which is the noisy feature we described in Fig. 12. The ensemble means are all smoother than the U-Net, expected from the spread of the ensemble members. Even the CorrDiff model (solid purple line) shows some blurriness overall compared to the observed image, but does have larger spectra for waves smaller than four km, compared to the U-Net. Considering the imagery in 12, the blurriness of the CorrDiff model manifests itself as softer edges to the area of developing convection. As for the LDM at 3 hours, there are some shorter wavelength information that matches the original image and some loss of waves at middle wavelengths, but the growth of artifacts near the edge is likely impacting the interpretation of the spectra. For future research we plan to use additional sharpness metrics to better quantify the impact of diffusion on overall sharpness of the machine learning output (sharpness2025).
@@ 4.2 Bulk Results
We compare here mainly to persistence and the U-Net forecast. Fig. 15 confirms what was shown qualitatively in Fig. 12 where most of the diffusion models have added value over the more traditional U-Net approach and a simple persistence forecast. Furthermore, all methods not only advect clouds (e.g., an optical flow solution; Shakya2019), they can also create and decay clouds. We are encouraged that these models work given the limited information content included as inputs.
@@
A key perk that we have seen already in the diffusion framework is the ability to generate n-many samples to use ensemble based forecasts. This already results in improved pixel-based metrics in Fig 15. We can also look to quantify how well the uncertainty of that ensemble corresponds to the true error of the model. We leverage the skill-spread diagram (e.g., Haynes2023; dellemonache2013probabilistic) where the standard deviation of the 10 member diffusion ensembles is compared to the RMSE of the ensemble (Fig. 16).
@@
5 Discussion
5.1 Lessons Learned
As seen in the prior subsections, developing diffusion models requires numerous design choices to navigate the trade-offs of practical applications. Here we summarize and discuss some of these choices and trade-offs:
  • We chose the overall framework of score-based diffusion models over DDPMs because they are significantly faster without presenting any disadvantages.
  • The underlying architecture, while arbitrary, does have an affect on the timeliness of the model. As we were learning to use diffusion we neglected to change the underlying architecture from an online tutorial. Turns out this architecture, while skillful, had many parameters and some attention layers that made training and inference times considerably large (5 days to train; 10 minutes to run a 3 hour forecast for 10 ensemble members). This resulted in us exploring latent diffusion models to try to speed up the model. An alternative route would have been to investigate simpler underlying architectures that are faster (e.g., less parameters, careful placement of attention layers), or explore other community provided and optimized networks like NVIDIA’s PhysicsNeMo which is used in Mardani2025 and Pathak2024.
  • Whether a latent model (LDM) should be used is a trade-off between higher inference speed and potentially lower accuracy. In our experience, conducting experiments with and without latent layers is the only way to tell whether they are worth using for a specific application. Experiments should include trying different latent space dimensions and different types of VAEs inside the LDM. A challenge is the large space of hyperparameters to explore in the LDM space, especially given that diffusion models are generally slow to train. Another challenge is training a VAE that faithfully reconstructs the training data. That is, any artifacts (blurring, added noise, token artifacts etc) created by the VAE are not correctable from the diffusion model and will end up in the final output.
  • An even bigger question for any specific application is whether diffusion models are worth the effort, including more complex code base and large computational costs for both training and inference. For example, ECMWF explored two options to extend AIFS - their AI-based global weather forecasting model (Lang2024a) - to provide probabilistic forecasts: (1) integrating diffusion models (Chantry2025AIFS), or (2) utilizing the Continuous Ranked Probability Score (CRPS) (Lang2024b) as a loss function to be used with a simpler neural network. After trying both approaches the CRPS route was chosen for first operationalization. Similarly, Google’s operational weather forecast models have seen a shift from Diffusion with their GenCast model (Price2025) to a similar CRPS loss with their FGN model (Alet2025), where the FGN model is now in operations. Following these two efforts, other global weather models have also used CRPS losses to get to ensemble forecasting (e.g., FuXi-ENS; Zhong2025), suggesting that those implementing weather forecasts have been focusing on implementation speed over potentially sharper forecasts with diffusion. Even though diffusion is currently costly in training and inference, there are many researching how to speed up diffusion (e.g., Xu2025).
5.2 Limitations and Future Directions